# Transaction Profiling and Address Role Inference in Tokenized U.S. Treasuries

### v1

@@ Abstract
Tokenized U.S. Treasuries have emerged as a prominent subclass of real-world assets (RWAs), offering cryptographically enforced, yield-bearing instruments collateralized by sovereign debt and deployed across multiple blockchain networks. While the market has expanded rapidly, empirical analyses of transaction-level behaviour remain limited. This paper conducts a quantitative, function-level dissection of U.S. Treasury-backed RWA tokens including BUIDL, BENJI, and USDY, across multi-chain: mostly Ethereum and Layer-2s. We analyze decoded contract calls to isolate core functional primitives such as issuance, redemption, transfer, and bridge activity, revealing segmentation in behaviour between institutional actors and retail users. To model address-level economic roles, we introduce a curvature-aware representation learning framework using Poincaré embeddings and liquidity-based graph features. Our method outperforms baseline models on our RWA Treasury dataset in role inference and generalizes to downstream tasks such as anomaly detection and wallet classification in broader blockchain transaction networks. These findings provide a structured understanding of functional heterogeneity and participant roles in tokenized Treasury in a transaction-level perspective, contributing new empirical evidence to the study of on-chain financialization.
keywords:
tokenized U.S. Treasuries, address role inference, transaction function analysis
category:
[…]
1 Introduction
@@
Real-world assets (RWAs) refer to off-chain, yield-bearing financial instruments such as sovereign debt, treasury, credit products, and real estate that are instantiated on-chain as cryptographically enforceable claims via tokenized representations [12], typically underpinned by compliance-bound issuance frameworks and asset custody infrastructures [21]. Early concept and implementation of RWA tokenization in decentralized finance (DeFi) pioneered in 2019 with the launch of Centrifuge’s Tinlake platform [22], which enabled the issuance of ERC-20 tokens [5] backed by tokenized real-world assets such as invoices and trade receivables. Subsequently, RWA tokens evolved to encompass a broader class of fixed-income and yield-generating instruments, notably short-duration U.S. Treasuries [18], tokenized real estate [24], private credit loans [3], each implemented through on-chain implementation of financial contracts governed by legal wrappers, custodial attestations, and programmable functions embedded in ERC-20 [5] or ERC-1400 [23] token standards.
Tokenized representations of short-duration U.S. Treasury bills have emerged as a dominant subclass within the RWA market, constituting over \$6.1 billion in on-chain assets as of Q2 2025 [20]. Leading instruments include BUIDL (BlackRock), OUSG (Ondo Finance), STBT (Matrixdock), and BENJI (Franklin Templeton) [15]. Economically, these tokens function as digital cash equivalents collateralized by Treasury securities, with daily yield accruals implemented via rebasing mechanics (e.g., rOUSG) [14] or dividend distribution (e.g., BUIDL) [25]. Smart contracts encode fund accounting logic, including NAV maintenance, dividend accrual, and investor disbursements, with compliance modules (e.g., Securitize) enforcing identity gating and transfer restrictions. The composability of these instruments within permissionless DeFi remains structurally limited due to non-transferability constraints and KYC-enforced whitelists. Despite this growing RWA U.S. treasury market, the on-chain behavioural patterns of these instruments across issuance, transfer, yield distribution, and institutional wallet interactions remain largely unexamined from a data-driven, motivating transaction-level analysis and network modeling of the emerging Treasury token market.
Prior empirical and smart contract-oriented research on RWA tokenization has centred on structural features of asset issuance and transfer mechanisms [4, 7], highlighting institutional governance trade-offs [28], composability limitations [7], and participation frictions [27, 11], while offering limited analysis of transactional behaviour from a data-driven perspective and cross-chain functional heterogeneity at scale. While those previous efforts were undertaken on RWA tokenization across various asset classes, focused empirical analysis of U.S. Treasury-backed tokens remains limited, even as these instruments have grown significantly since recent years, with increasing contract deployments and transaction data generated flows spanning various financial functionalities.
To address persistent analytic gaps in the study of tokenized real-world assets, this work conducts a function-decoded, cross-chain dissection of tokenized U.S. Treasury transaction networks, isolating transaction-level functional primitives: issuance, redemption, transfer, and bridging, to delineate institutional versus retail financial behaviours. This multi-chain functional analysis enables the differentiation of institutional and retail usage patterns by comparing transaction frequency, notional value, and function-specific activity both within individual chains and in aggregate across the RWA token protocols. The analysis is grounded in transaction-level data collected across multiple chains for three representative U.S. Treasury-backed RWA tokens: BUIDL (BlackRock USD Institutional Digital Liquidity Fund), BENJI (Franklin OnChain U.S. Government Money Fund), and USDY (Ondo U.S. Dollar Yield). We further propose a role classification model for addresses by integrating hyperbolic representation learning, liquidity-based behavioural features, and hierarchical graph descriptors to infer latent economic functions such as treasuries bots, and retail traders.
1.1 Overview
To the best of our knowledge, this is the first study to systematically collect, decode, and structure transaction-level data for tokenized U.S. Treasuries RWA tokens, supporting a quantitative, function-level analysis of these assets from a data-driven perspective. Our contributions are as follows:
  • Cross-chain functional decomposition We present a contract-decoded, function-level analysis of tokenized U.S. Treasuries across multiple chains, revealing how issuance, redemption, transfer, and DeFi-related operations vary by chain and token design.
@@ 1.1 Overview
  • Investor type via transactional profiling: We perform statistical segmentation of wallet behaviour using transaction frequency and value distributions, identifying differences between institutional-scale actors and retail participants. Our results show that institutional wallets dominate issuance and redemption flows, while retail users engage primarily in mid-value transfer activity concentrated on Layer-2 networks.
@@
  • Address role inference via curvature-aware representation learning: We propose a predictive framework based on Poincaré embeddings to capture latent transaction geometry for address role inference (e.g., treasury, bots, traders). Our approach outperforms established baselines on the RWA Treasury dataset and exhibits competitive generalization to downstream tasks such as anomaly detection and entity classification on diverse public blockchain transaction network datasets.
2 Related Work
@@
Real-world asset (RWA) tokenization has increasingly focused on fixed-income instruments, with recent studies examining blockchain-based representations of sovereign debt, particularly tokenized U.S. Treasuries, as a representative use case within on-chain capital markets [4]. A recent multi-sector case study similarly reports that tokenization can improve transaction efficiency and create new value, yet it also introduces governance complexities and shifts risk distribution in on-chain markets [28]. In addition, empirical surveys of thirty-nine RWA projects have been studied to reveal common on-chain vulnerabilities, such as heavy reliance on stablecoins for settlement and a limited base of active on-chain investors [7].
@@
From a behavioural perspective, research efforts have been undertaken examine address-level data to characterize RWA usage patterns. Swinkels studied Ethereum-based real estate tokens (fractional shares of 58 U.S. rental properties) and found highly fragmented ownership, with around 254 unique holders per property on average [27]. The author observed that larger token holders tended to spread investments across multiple properties, while overall liquidity was low: each property token changed ownership only about once per year (slightly more often if tradable on decentralized exchanges). Kreppmeier et al. tracked 173 U.S. real estate security tokens (over 238,000 on-chain transactions) and similarly found broad participation by small investors, though individual wallets were not well-diversified across different tokens [11]. In addition to property fundamentals, the study showed that crypto-market factors such as transaction costs and investor sentiment significantly influence both initial token offering success and subsequent trading flows. These studies demonstrate the value of categorizing on-chain addresses by their roles and behaviours, for instance, distinguishing issuers, custodians, and various types of investors to better interpret RWA economical patterns. The emerging of U.S. Treasury RWA tokens, i.e., on-chain shares in Treasury-backed funds, has drawn interest for bringing safe assets on-chain [18], although academic analysis of their transaction-level semantics such as participant composition remains scant.
3 Data Collection
@@
We collect raw on-chain transactions for tokenized real-world asset products representing U.S. Treasuries, namely, BUIDL (BlackRock USD Institutional Digital Liquidity Fund), BENJI (Franklin OnChain U.S. Government Money Fund), and USDY (Ondo U.S. Dollar Yield) across multiple blockchains using public blockchain explorer APIs. Our collection pipeline queries EVM-compatible chains (e.g., Ethereum, Arbitrum, Mantle), as well as non-EVM chains (e.g., Aptos, Stellar), and fetches transfers for the three most prominent RWA tokens by market capitalization: BUIDL, BENJI, and USDY, based on rankings from market cap  [^1] https://app.rwa.xyz/treasuries.
@@
Table 1: Cross-chain summary statistics for tokenized U.S. Treasury instruments. Columns report the number of transactions, unique participating addresses, and the range of observed activity.
@@
BUIDL shows the most activity on Ethereum, aligning with its institutional adoption, though it is collected across multiple chains. The BENJI token exhibits exceptionally high activity on Stellar, with over 2.5 million transactions involving more than 220,000 addresses in just two weeks, indicating high-frequency issuance or custodial activity. USDY shows dense activities on Mantle, accounting for more than 274,000 transactions. Overall, this collected multi-chain dataset enables us to quantify RWA adoption patterns and behavioural clusters in later sections. To decode smart contract method calls, we utilize Tenderly’s decoding engine  [^2] https://tenderly.co/ and the open-source 4byte signature database  [^3] https://www.4byte.directory/ to decode the input field of each transaction, enabling function-level analysis. The result is a cross-chain, function decoded dataset of RWA token transfers suitable for downstream quantitative analysis.
4 Quantitative Analysis of RWA Tokens
To quantitatively assess whether U.S. Treasury token transactions align more closely with institutional investors or retail users, we analyze the frequency of transactions versus their size. An institutional usage pattern is characterized by relatively few transactions, each of very large notional value, while a retail-driven pattern involves a high frequency of smaller transactions, typically ranging from a few dollars to several thousand. We further stratify this analysis by both chains and functions, to assess how usage patterns vary across blockchain environments (e.g., Ethereum vs. Arbitrum) and operational roles (e.g., issuance, redemption, transfer). This enables us to disentangle how token delivers differently in function-level interaction and chain-specific deployment.
Figure 1: Log-scaled scatter plot of BUIDL token interactions by function and chain. Each point represents a distinct (function, chain) pair, where the x-axis is the total BUIDL value transacted and the y-axis is the transaction count. Marker shape denotes functional class (e.g., transfer, issuetokens, redeem), and colour indicates blockchain deployment (e.g., Ethereum, Arbitrum). Institutional behaviour is typically clustered at the top-right (few, large transactions), while retail-aligned activity appears bottom-left (many, small transactions).
4.1 BUIDL
Table 2: Functional buckets for decoded contract functions. Substrings are matched against decoded function names. Function reflect high-level operational roles in RWA token flows.
Function buckets Name Match Description
issuetokens issue, mint, bulkissuance Initial minting and dividend payments to investors
@@ 4.1 BUIDL
transfer transfer, bridgedstokens, multisend Routine transfers or bridging between chains
@@
We categorize smart contract decoded functions into operational buckets based on string-matching heuristics against lowercase function names to interpret the functional semantics. Table 2 summarizes these mappings, which allow for cross-chain, token-agnostic aggregation of function activity into economically meaningful categories. For instance, the issuetokens bucket includes all minting operations ranging from initial token creation to dividend payment (e.g., via batched bulkIssuance calls in BUIDL). Likewise, redeem captures capital outflows by investors, while burn reflects permanent supply contraction. The transfer bucket encompasses routine value transfers, including bridged hops and multi-send operations.
@@
Figure 1 visualizes the distribution of BUIDL token interactions by function and chain, where each data point corresponds to a unique $(f,c)$ pair: a specific function (bucket) $f$ executed on chain $c$. The x-axis (log scale) captures the total value transferred through the function, while the y-axis captures the number of transactions. Marker shapes denote function classes (e.g., transfer, issuetokens), and colours distinguish the underlying chains (e.g., Ethereum, Polygon, Arbitrum). Points located in the upper-right quadrant represent functions with both high frequency and high cumulative value. In contrast, points near the upper-left indicate high-frequency, low-value interactions consistent with retail-like activity, and points in the lower-right denote low-frequency but high-value functions, as expected for capital-intensive institutional transfers.
@@
The observed distribution of BUIDL’s on-chain activity is heavily skewed toward the lower-right, especially on Ethereum, with large-value, low-frequency function calls such as redeem, burn, and issuetokens dominating the volume. This pattern is consistent with usage by accredited or institutional actors, rather than mass-market retail adoption. The dominance of Ethereum reinforces its role as the canonical execution environment for regulated, institutional-grade RWA issuance and settlement. However, the presence of BUIDL transactions on alternative chains such as Avalanche, Polygon, and Optimism indicates emerging interoperability requirements. These deployments may support secondary custody flows, bridging infrastructure, or protocol-level integrations where gas efficiency or modular composability is prioritized. Such cross-chain activity for BUIDL remains limited as of April 2025, based on our collected data, with the vast majority of transaction volume and functional engagement concentrated on Ethereum. Notably, BUIDL transactions are almost exclusively initiation-side operations (e.g., primary issuance and redemption), with minimal secondary trading activity, suggesting that most interacting addresses represent institutional counterparties rather than retail participants.
4.2 BENJI
@@
Franklin Templeton’s BENJI token representing shares of its OnChain U.S. Government Money Fund, was initially deployed on the Stellar blockchain in 2021 and has since used Stellar as the primary ledger for recording share ownership and transactions [26]. In 2023–2024, BENJI was expanded beyond Stellar onto several EVM-compatible chains: Polygon, Arbitrum, Avalanche, and Ethereum Mainnet. Each blockchain in the BENJI token’s plays a specialized role, forming an integrated cross-chain architecture rather than independent silos. Stellar still functions as the primary issuance and settlement network, where the majority of BENJI tokens are minted and held (most of the token supply was on Stellar as of April 2024 [9]).
@@
Table 3: Transaction summary for BENJI across supported chains and decoded functions, based on data collected through April 2025. Stellar shows exceptionally high activity due to its role as the fund’s primary settlement ledger, while EVM chains primarily use the signedDataExecution function for controlled, authorized operations.
@@
Chain Function Transanction Total Value First Seen Last Seen
@@
4.3 USDY
We also categorize USDY smart contract functions into operational buckets by applying heuristic string matching on the lowercase decoded function names, aiming to infer their functional semantics. Table 4 describes this mapping, specifically for the USDY token’s diverse DeFi interactions. Given the wide variety of protocols and contract architectures interacting with USDY across multiple chains, the mapping produced aggregation based on the functional semantics. For example, swap captures all decentralized exchange operations regardless of DEX type, while lending encompasses a range of credit-related actions such as borrow, repay, or collateral management.
@@ 4.3 USDY
Table 4: Functional buckets for USDY contract activity. Buckets are derived by substring pattern matching in decoded function names. These categories reflect common DeFi roles in cross-chain real-world asset token ecosystems.
Function Bucket Name Match (substring) Description
@@
bridge bridge, startbridge, swapandstartbridge Cross-chain bridging of tokens
@@
execution executemeta, execute, exectransaction, delegatecall, call, multicall General-purpose call wrappers and execution shells
@@
Figure 2 visualizes the cross-chain function-level distribution of USDY transactions by plotting total token value against transaction count for each (function, chain) pair. The on-chain data for USDY reveals a mix of institutional-sized wallets and numerous retail holders. The function-level transaction data for USDY reveals a dual-structure usage pattern indicative of both institutional and retail engagement across multiple blockchain environments. Despite USDY’s positioning as a regulated, yield-bearing stablecoin offered under Regulation S [29] to non-U.S. investors, on-chain evidence suggests that the asset has achieved a relatively broad distribution: over 10,000 unique addresses hold USDY across supported chains, suggesting significant uptake by smaller-scale retail participants. However, transaction size and function type diverge markedly between user classes. Large-value, low-frequency operations—such as mint, burn, and high-value execution calls are primarily concentrated on Ethereum mainnet, consistent with institutional-scale issuance and redemption flows. These interactions exhibit transaction sizes in the hundreds of thousands to millions of USDY, aligning with primary market activity and custody-level fund management. By contrast, Layer-2 networks such as Arbitrum and Mantle exhibit distinct transactional profiles characterized by higher frequency, lower median value, and increased heterogeneity in function types (e.g., swap, transfer, and permissioned execution). These patterns are congruent with DeFi-native retail usage, enabled by the lower transaction costs and faster finality on L2s. The structural constraint, though conducive to regulatory compliance, inherently limits composability with standard ERC-20 interfaces and precludes fully permissionless usage; further bifurcating institutional administrative flows from retail DeFi interactions across different layers of the chain stack.
5 Address Role Predictive Modeling

### v2

# Decoding RWA Tokenized U.S. Treasuries: Functional Dissection and Address Role Inference

@@ Abstract
Tokenized U.S. Treasuries have emerged as a prominent subclass of real-world assets (RWAs), offering cryptographically secured, yield-bearing instruments issued across multi-chain Web3 infrastructures, with growing significance for transparency, accessibility, and financial inclusion. While the market has expanded rapidly, empirical analyses of transaction-level behaviours remain limited. This paper conducts a quantitative, function-level dissection of U.S. Treasury-backed RWA tokens, including BUIDL, BENJI, and USDY across multi-chain: mostly Ethereum and Layer-2s. Decoded contract calls expose core financial primitives such as issuance, redemption, transfer, and bridging, revealing patterns that distinguish institutional participants from smaller or retail users for the extent and limits of inclusivity in current RWA adoption. To infer address-level economic roles, we introduce a curvature-aware representation learning model. Our method outperforms baseline models in role inference on our collected U.S. Treasury transaction dataset and generalizes to address classification across broader public blockchain transaction datasets. The decoded transaction-level patterns in tokenized U.S. Treasuries across chains surface the degree of retail participation, and the role inference model enables the distinction between institutional treasuries, arbitrage bots, and retail traders based on behavioral patterns, facilitating future more transparent, inclusive, and accountable Web3 finance.
I Introduction
@@
Real-world assets (RWAs) refer to yield-bearing financial instruments such as sovereign debt, treasuries, credit products, and real estate that are instantiated on-chain as cryptographically enforceable claims via tokenized representations [24], typically underpinned by compliance-bound issuance frameworks and asset custody infrastructures [41]. Subsequently, RWA tokens have evolved to encompass a broader class of fixed-income and yield-generating instruments, notably short-duration U.S. Treasuries [37], tokenized real estate [43], private credit loans [5], each implemented through on-chain financial contracts governed by legal wrappers, custodial attestations, and programmable functions embedded in ERC-20 [9] or ERC-1400 [42] token standards. Such tokenized assets are increasingly viewed through the lens of financial inclusion, particularly for enabling access to sovereign debt instruments across geopolitical and socioeconomic boundaries, e.g., allowing individuals in emerging markets or lacking U.S. brokerage accounts to access U.S. Treasuries via on-chain platforms.
Tokenized representations of short-duration U.S. Treasury bills have emerged as a dominant subclass, expected to reduce issuance or settlement frictions and compliance costs, as well as broaden global (including retail) access to mainstream securities [12]. Such instruments complement DeFi as safe collateral and programmable primitives [4], constituting over \$6.1 billion in on-chain assets as of Q2 2025 [40]. Among leading instruments, we focus on BUIDL (BlackRock), BENJI (Franklin Templeton), and USDY (Ondo): the largest by market capitalization and on-chain activity as of Q2 2025 [39]. Functionally, these tokens implement yield accruals via rebasing mechanics [33] or dividend distribution [44], and embed smart contracts that encode the logic and enforce verifiable transfers. Despite the market’s growth, it remains unclear whether such instruments meaningfully expand access beyond institutional actors to retail participants. Moreover, the transparency and role of on-chain behavior in shaping an inclusive and accountable financial infrastructure remain insufficiently understood.
Prior research on RWA tokens centred only on features of asset issuance and transfer mechanisms [8, 13], institutional governance trade-offs [47], composability limitations [13], and participation frictions [46, 23]. To date, no studies have extracted transaction-level data to construct function-decoded, cross-chain token transfers and conduct address role modeling, the potentially foundational components for analyzing whether RWA tokenization benefits more equitable access financial systems, and broader socioeconomic inclusion in line with Sustainable Development Goals [7].
Concretely, three critical deficiencies: The initial challenge is dataset construction: no public resource has aggregated a cleaned, formatted collection of an RWA U.S. Treasury transaction dataset, with normalized chain-specific metadata, and supplied tractable address (role) labels across multiple blockchains. The absence of such a processed, multi-chain dataset precludes reproducible cross-chain analytics and limits empirical validation of claims. Secondly, no prior study has undertaken a transaction-level functional audit of RWA U.S. Treasury token contracts: The transactions have not been decomposed into economic primitives such as issuance, redemption, liquidity provision, bridging; nor have the analyzed results that have been traced to insights of research questions including (i) quantify institutional versus retail participation; which group exhibits broader engagement in the market. (ii) characterize how individual chains specialise within the multi-blockchain deployment stack, clarifying why specific chain hosts some distinct financial functionalities. In addition, the literature lacks a principled weak-supervision model capable of mapping on-chain addresses to behavioral roles (e.g., custodian, treasury, execution bot, trader), thereby limiting downstream applications such as compliance screening [50], AML surveillance [2], and liquidity-risk attribution [14]. Such modeling also strengthens transparency and accountability in blockchain-based financial systems, advancing infrastructure aligned with goals related to inclusive economic growth and reduced inequalities [7]. To our knowledge, this is the first study to systematically structure transaction-level data for tokenized U.S. Treasuries, enabling quantitative, function-level analysis. Our contributions are as follows:
U.S. Treasury Tokens Transaction Dataset: We assemble a multi-chain transaction dataset for RWA U.S. Treasury tokens, including BUIDL, BENJI, and USDY, mapping verified contracts, decoding method calls, and collecting community-accumulated address annotations with simplified role labels to enable reproducible, function-level analysis; the dataset will be made available upon publication.
@@
Cross-chain functional decomposition: We present a contract-decoded, function-level analysis of tokenized U.S. Treasuries across multiple chains, revealing how issuance, redemption, transfer, and DeFi-related operations vary by chain and token design. We perform statistical segmentation of wallet behaviour using transaction frequency and value distributions, identifying differences between institutional-scale actors and retail participants. Our results show that institutional wallets dominate issuance and redemption flows, while retail users engage primarily in mid-value transfer activity concentrated on Layer-2 networks.
@@
Address role inference model: We propose a predictive model based on Poincaré embeddings to capture latent transaction geometry for address role inference (e.g., treasury, bots, traders). Our approach outperforms established baselines on the RWA Treasury dataset and exhibits competitive generalization to downstream tasks such as anomalous address classification on diverse public blockchain transaction network datasets.
II Related Work
@@
Real-world asset tokenization has increasingly focused on fixed-income instruments, with recent studies examining blockchain-based representations of sovereign debt, particularly tokenized U.S. Treasuries, as a representative use case within on-chain capital markets [8]. A recent multi-sector case study similarly reports that tokenization can improve transaction efficiency and create new value, yet it also introduces governance complexities and shifts risk distribution in on-chain markets [47]. In addition, empirical surveys of thirty-nine RWA projects have been studied to reveal common on-chain vulnerabilities, such as heavy reliance on stablecoins for settlement and a limited base of active on-chain investors [13].
@@
From a behavioural perspective, research efforts have been undertaken to examine address-level data to characterize RWA usage patterns. Swinkels studied Ethereum-based real estate tokens (fractional shares of 58 U.S. rental properties) and found highly fragmented ownership, with around 254 unique holders per property on average [46]. The author observed that larger token holders tended to spread investments across multiple properties, while overall liquidity was low: each property token changed ownership only about once per year (slightly more often if tradable on decentralized exchanges). Kreppmeier et al. tracked 173 U.S. real estate security tokens (over 238,000 on-chain transactions) and similarly found broad participation by small investors, though individual wallets were not well-diversified across different tokens [23]. In addition to property fundamentals, the study showed that crypto-market factors such as transaction costs and investor sentiment significantly influence both initial token offering success and subsequent trading flows. These studies demonstrate the value of categorizing on-chain addresses by their roles and behaviours, for instance, distinguishing issuers, custodians, and various types of investors to better interpret RWA economical patterns. The emerging of U.S. Treasury RWA tokens, i.e., on-chain shares in Treasury-backed funds, has drawn interest for bringing safe assets on-chain [37], although academic analysis of their transaction-level semantics such as participant composition remains scant.
III U.S. Treasury Tokens Dataset Collection
@@
We collect raw on-chain transactions for tokenized real-world asset products representing U.S. Treasuries, namely BUIDL (BlackRock USD Institutional Digital Liquidity Fund), BENJI (Franklin OnChain U.S. Government Money Fund), and USDY (Ondo U.S. Dollar Yield) across multiple blockchains using public blockchain explorer APIs. Our collection pipeline queries EVM-compatible chains (e.g., Ethereum, Arbitrum, Mantle), as well as non-EVM chains (e.g., Aptos, Stellar), and fetches transfers for the three most prominent RWA tokens by market capitalization: BUIDL, BENJI, and USDY based on rankings from market cap [39]. We enumerate and resolve all verified smart contract deployments across their supported chains, extracting token transfer activity via the following official block explorers: Etherscan (Ethereum), Polygonscan (Polygon), Arbiscan (Arbitrum), Basescan (Base), Optimistic Etherscan (Optimism), Snowtrace (Avalanche), and Aptoscan (Aptos). Transaction logs were retrieved from contract inception through April 2025, spanning 11 chain-token pairs.
@@
TABLE I: Cross-chain summary statistics for tokenized U.S. Treasury. Columns report the number of transactions, unique participating addresses, and the range of observed activity.
@@
BUIDL shows the most activity on Ethereum, aligning with its institutional adoption, though it is collected across multiple chains. The BENJI token exhibits exceptionally high activity on Stellar, with over 2.5 million transactions involving more than 220,000 addresses in just two weeks, indicating high-frequency issuance or custodial activity. USDY shows dense activities on Mantle, accounting for more than 274,000 transactions. Overall, this collected multi-chain dataset enables us to quantify RWA adoption patterns and behavioural clusters in later sections. To decode smart contract method calls, we utilize Tenderly’s decoding engine [48] and the open-source 4byte signature database [17] to decode the input field of each transaction, enabling function-level analysis. The result is a cross-chain, function-decoded dataset of RWA token transfers suitable for downstream quantitative analysis.
IV Quantitative Analysis of RWA Tokens
To quantitatively assess whether U.S. Treasury token transactions align more closely with institutional investors or retail users, we analyze the frequency of transactions versus their size. An institutional usage pattern is characterized by relatively few transactions, each of very large notional value, while a retail-driven pattern involves a high frequency of smaller transactions, typically ranging from a few dollars to several thousand. We further stratify this analysis by both chains and functions to assess how usage patterns vary across blockchain environments (e.g., Ethereum vs. Arbitrum) and operational roles (e.g., issuance, redemption, transfer).
Figure 1: Log-scaled scatter plot of BUIDL token transactions by function and chain. Each point represents a (function, chain) pair, where the x-axis is the total BUIDL value transacted and the y-axis is the transaction count. Marker shape denotes functional class (e.g., transfer, issuetokens, redeem), and colour indicates a blockchain (e.g., Ethereum, Arbitrum).
TABLE II: Functional buckets for decoded contract functions. Substrings are matched against decoded function names. Functions reflect high-level operations in RWA token flows.
Func buckets Name Match Description
issuetokens issue, mint, bulkissuance Initial minting and dividend payments
transfer
transfer, bridgedstokens,
multisend
Routine transfers or bridging
between chains
@@ IV Quantitative Analysis of RWA Tokens
BUIDL. We categorize the contract decoded functions into operational buckets based on string-matching heuristics against function names to interpret the functional semantics. Table II summarizes these mappings, which allow for cross-chain, token-agnostic aggregation of function activity into economically meaningful categories. For instance, the issuetokens bucket includes all minting operations ranging from initial token creation to dividend payments (e.g., via batched bulkIssuance calls in BUIDL). Likewise, redeem captures capital outflows by investors, while burn reflects permanent supply contraction. The transfer bucket encompasses routine value transfers, including bridged hops and multi-send operations.
@@
Figure 1 visualizes the distribution of BUIDL token interactions by function and chain, where each data point corresponds to a unique $(f,c)$ pair: a specific function (bucket) $f$ executed on chain $c$. The x-axis (log scale) captures the total value transferred through the function, while the y-axis captures the number of transactions. Marker shapes denote function classes (e.g., transfer), and colours distinguish the underlying chains (e.g., Ethereum, Polygon, Arbitrum). Points located in the upper-right quadrant represent functions with both high frequency and high cumulative value. In contrast, points near the upper-left indicate high-frequency, low-value interactions consistent with retail-like activity, and points in the lower-right denote low-frequency but high-value functions, as expected for capital-intensive institutional transfers.
@@
The observed distribution of BUIDL’s on-chain activity is heavily skewed toward the right, especially on Ethereum, with large-value function calls such as redeem, transfer, burn, and issuetokens dominating the volume. This pattern is consistent with usage by accredited or institutional actors, rather than mass-market retail adoption. The dominance of Ethereum reinforces its role as the canonical execution environment for regulated, institutional-grade RWA issuance and settlement. However, the presence of BUIDL transactions on alternative chains such as Avalanche, Polygon, and Optimism indicates emerging interoperability requirements. These deployments may support secondary custody flows, bridging infrastructure, or protocol-level integrations where gas efficiency or modular composability is prioritized. Such cross-chain activity for BUIDL remains limited as of April 2025, based on our collected data, with the vast majority of transaction volume and functional engagement concentrated on Ethereum. Notably, BUIDL transactions are almost exclusively initiation-side operations (e.g., primary issuance and redemption), with minimal secondary trading activity, suggesting that most interacting addresses represent institutional counterparties rather than retail participants.
@@
BENJI. Franklin Templeton’s BENJI token representing shares of its OnChain U.S. Government Money Fund, was initially deployed on the Stellar blockchain in 2021 and has since used Stellar as the primary ledger for recording share ownership and transactions [45]. From 2023 to 2024, BENJI was expanded beyond Stellar onto several EVM-compatible chains: Polygon, Arbitrum, Avalanche, and Ethereum. Each blockchain of the BENJI token’s plays a specialized role. Stellar still functions as the primary issuance and settlement network, where the majority of BENJI tokens are minted and held (most of the token supply was on Stellar as of April 2024 [20]).
@@
TABLE III: Transaction summary for BENJI across supported chains and decoded functions, based on data collected through April 2025. Stellar shows exceptionally high activity due to its role as the fund’s primary settlement ledger, while EVM chains primarily use the signedDataExecution function for controlled, authorized operations.
@@
Chain Function Transaction Total Value First Seen Last Seen
@@
USDY. We also categorize USDY smart contract functions into operational buckets by applying heuristic string matching on the decoded function names. Table IV describes this mapping. Given the variety of protocols and contracts interacting with USDY across multiple chains, the mapping produced an aggregation based on the functional semantics. For example, swap captures all decentralized exchange operations regardless of type, while lending encompasses credit-related actions such as borrow, repay, or collateral management.
@@
TABLE IV: Functional buckets for USDY contract activity. Buckets are derived by substring pattern matching in decoded function names. These categories reflect common DeFi roles in cross-chain real-world asset tokens.
Func Bucket Name Match (substring) Description
@@
bridge
bridge, startbridge,
swapandstartbridge
Cross-chain bridging of tokens
@@
execution
executemeta, execute,
exectransaction,
delegatecall, call, multicall
General-purpose call wrappers and
execution shells
@@
Figure 2 visualizes the cross-chain function-level distribution of USDY transactions by plotting total token value against transaction count for each (function, chain) pair. The on-chain data for USDY reveals a mix of institutional-sized wallets and numerous retail holders. The function-level transaction data for USDY reveals a pattern indicative of both institutional and retail engagement across multiple blockchains. Despite USDY’s positioning as a regulated, yield-bearing stablecoin offered under Regulation S [36] to non-U.S. investors, on-chain evidence suggests that the asset has achieved a relatively broad distribution: over 10,000 unique addresses hold USDY across supported chains, suggesting significant uptake by smaller-scale retail participants. However, transaction size and function type diverge markedly between user classes. Large-value, low-frequency operations such as mint, burn, and high-value execution calls are primarily concentrated on Ethereum, consistent with institutional-scale issuance and redemption flows. These interactions exhibit transaction sizes in the hundreds of thousands to millions of USDY, aligning with primary market activity and custody-level fund management. By contrast, Layer-2 networks such as Arbitrum and Mantle exhibit distinct transactional profiles characterized by higher frequency, lower median value, and increased heterogeneity in function types (e.g., swap, transfer, and permissioned execution). These patterns are congruent with DeFi-native retail usage, enabled by the lower transaction costs and faster finality on L2s. The structural constraint, though conducive to regulatory compliance, limits composability with standard ERC-20 interfaces and precludes permissionless usage; further bifurcating institutional administrative flows from retail DeFi interactions across different layers of the chain stack.
V Address Role Predictive Modeling